-The setup procedure takes 1-2hrs to complete once everything is powered up.
-In summer start as late as possible to avoid alignment problems resulting 
 from solar heating of the tunnel/enclosure.
-You might need to check the internal alignment after a few (3-4?) hrs as 
 tunnel temperature cools.
-Shutdown and logout each night otherwise a server lag settles in and
 causes trouble.

-Power up the building and systems if noone has been 
 at SUSI for a while.

-Check vacuum pipe pressure. If higher than 0.2atm pump down for ~1hr to ~0.05atm.
-Select your sids for the night, depending on observing program.
-Login as Observer, pwd is on noticeboard.
-Open 4 terminals: Login to Peleas(ssh), Gareth(ssh+passwd), Gaheris(ssh+passwd).
 Leave one as Arthur
-Check all clocks match, 'date'.
-Run picoset as observer on arthur. This sets all the picomotor speeds.
-Start up UIs:
  In Peleas:  xrun_sidcon ?x    (?=n/s, x=1-4)
  In Gareth:  xrun_fringecon
  In Gaheris: xrun_wobble2
              wobble2gtk (do this after completing the wobble2 autolist)
                         (wobble2gtk should work from any machine)
  In arthur:  sidcongtk ?x  (where ?x matches the sidcons above)
                (Do this after the sidcon servers have completed their autolist)
              xrun_scheduler & [Starts xmms (sounds) and ccat (command history)]

-For each UI follow the "automatic setup procedure/run the auto list", 1

-Sidcons: Put periscopes in
-LDC: Ensure LEDS are on. Otherwise the limit switches arent turned on and the wedge
      runs into the endstops and shreds the pulley!
      Check Mode is 'Air & Glass', wavelength, and that it indexes correctly.
      Now set to zero total optical dispersion (2 5 ie. from main menu go to 2 then 5). 
      May need to hit red LDC 'reset' button on north rack.
-Wobble2: check camera temp & sink (~-40to-50, ~27to32deg is normal)
-Opttab2: srcin,  rdsout,
          inpin,  assout,
          refin,  edsin,
          selout, nmsout,
          monout, smsout,
-fringecon_client: server|connect
  Check settings in fringecon (LUT, sample time, filter, damping, process type:
  eg. scan140_1024(long) or scan35f_256(short), 0.2, 0.7/0.08, 4, equalise apds)
-opcon: You may need to hit the red PLC 'reset' button on north rack if it
         doesn't start.
        type 'centre' to turn both piezos on and centre them
        [OR from main menu:2 2 2 1 select piezo and turn both on (4 to centre->0x800)]
        from main     :6 to toggle status display
        from main:    :4 for tracking menu
        If the dogleg is in type 'user_offset n -72106953'

        Zero the carriage in PLC
        2 1
        5 to centre piezos
        3 2 1 set rate|unstop|stop to move carriage to zero position
              use rates 10,5,2 near centre and end by going +ve
        0  load centre position values, ie. zero it
        (You will need to unplug/replug the zeroset guage box)
-Start StarTrackMsg (stMsg) on PC. Must do this AFTER starting wobble2 & wobble2gtk
   (Note this startup order for times when wobble2 is shutdown/restarted)

-Scheduler should make and set a new data directory (msd) on opening
  Do  dir /export/snert/data/yymmdd (if necessary)
        if this fails type msd in an arthur window
      sid n? s? ?     to set the sids an direction, eg. sid n4 s3 s
      reopen          to make all connections 
      fs 300 600 15   eg. to set fringe search parameters 
      Note: scans 1000 50 is default - see Scheduler notes.
-wobble2gtk: Take darks in track, ref and widefield modes to check response,
             ie. that the counts drop to ~zero. Lights in enclosure should be off.

-Ensure stMsg intensifier (black box under monitor), APDs and CCD are OFF
-Connect ports #54 and #55 with the short patch cable. This connects the
 Pico-motor Hand Paddle (PHP) to port #55 by the BRT.
-Into enclosure. Take PHP. Take a siderostat key. Switch on laser, remove all ND. 
 Leave polariser in. If necessary, adjust src mirror to centre laser on pinhole,
 ie. to ensure even illumination of pupil.
-Check small apertures (10mm) are in. Check laser spots appear on BRT double target.
-Switch on wobbler power supplies (near BRT). They should settle at 72.5V.

-Go to sids, remove bins and heaters, unbolt roof. Leave rooves closed till dark.

And finally,
-Back to BRT, follow alignment procedure

-Stop tracking in sids. Stow sids (ensure they stop at El~330deg)
-Stop tracking in opcon, slew to zero. Zero up.
-LDC:2 5 to set to zero dispersion position
-Wobble2gtk: open servos, stop logging data, stop.
-To electronics rack: sids and OPLC to standby, air-con on.
-Exit stMsg, turn off intensifier
-Close down all servers
-Switch off APDs. Check intensifier is off and CCD is stopped.
-Into enclosure: ND out of laser, check all is OK
-Turn off wobbler power supplies
-Put BRT double target on. Put BRT lower periscope dust covers on.
-To sids: bin the neds, mirror heater in place, close housings
-Complete log
-Log out
-Go to bed and sleep:)
Powering up the building and systems
First do the control room then go clockwise around the tunnel.
Reverse this list to power down.

Control Room
-South (red) electronics rack - at back
 -switch on 3(4?)x GPOs at bottom of middle rack marked 'red table'
 -switch on red-table power supplies box (red switch)
-North (blue) electronics rack - at back
 -NOTE: on shutdown leave power on to panel holding air-con switches.
        This will allow local staff to switch a/c on before we arrive.
 -switch on 8x GPOs on left-hand rack, all labelled
 -check all power leads (both ends) are well seated
 -ensure a fan is blowing over the LDC box to prevent frequent
  LDC RTP failures
-Dynavac vacuum pipe gauge in NE corner
-Printer, photocopier and fax, air-con in control room and living room

 Going clockwise from NE cnr of control room
-Blue table tunnel
 -1(2?)x GPO. Supplies Pico 8,9
-North end of tunnel
 -UPS.Check 4x GPO on enclosure roof. Supplies south metrology and Pico 10,11
-Centre of OPLC (inside enclosure)
 -UPS under OPLC near wobblers. Supplies carriage
-BRT tunnel on enclosure roof
 -3x GPO -UPS.Supplies acquisition camera + other
         -Environdata. LEAVE ON.
         -Pico 4,5,6,7
-South tunnel (near BRT tunnel)
 -UPS. Supplies wobbler power supplies
-South end of tunnel (outside enclosure, western side)
 -UPS. Supplies north metrology
-South end of tunnel (inside enclosure)
 -1x GPO. Supplies Pico 11,12
-Red table tunnel
 -UPS. Supplies power board A. LEAVE ON.
  -Power board A -CCD cooler,  LEAVE ON unless leaving for a long time.
                 -CCD          (Note order: For on, Cooler then CCD and vice versa)
 		 -APD + scanning mirror
 		 -Pico 1,2,3 + Pico ethernet controller (must be switched on
		  *after* all pico motors have been switched on. Note you must
		  also reset all the motor speeds whenever the controller has
		  been switched off with 'picoset')
-Red table
 -GPO on floor at south side of red table. 
  Supplies alignment laser, white light source 
  and filter wheels (leave fw's off usually)


Andrew Jacob
Last modified: Wed May 5 05:26:35 EST 2004